If a computer program draws aspiral at a reate of 45% every 20 seconds then how long will it take to complete 25 full turns?
Answer by checkley77(12844)

I ASSUME THE 45% SHOULD BE 45 DEGREES.
25*360/45*20=25*8*45=9,000 SECONDS.